a rope is drawn around a circular land if the length of rope is 220m what is the radius
Answer by stanbon(75887)

a rope is drawn around a circular land if the length of rope is 220m what is the radius
-----
Circumference = 2pi*r
---
Equation:
220m = 2*pi*r
-----
radius = 220/(2pi) = 35.01 m
===================
Cheers,
Stan H.